### v1

# Enhanced-FQL($\lambda$), an Efficient and Interpretable RL with novel Fuzzy Eligibility Traces and Segmented Experience Replay

Abstract
This paper introduces a fuzzy reinforcement learning framework, Enhanced-FQL($\lambda$), that integrates novel Fuzzified Eligibility Traces (FET) and Segmented Experience Replay (SER) into fuzzy Q-learning with Fuzzified Bellman Equation (FBE) for continuous control tasks. The proposed approach employs an interpretable fuzzy rule base instead of complex neural architectures, while maintaining competitive performance through two key innovations: a fuzzified Bellman equation with eligibility traces for stable multi-step credit assignment, and a memory-efficient segment-based experience replay mechanism for enhanced sample efficiency. Theoretical analysis proves the proposed method convergence under standard assumptions. Extensive evaluations in continuous control domains demonstrate that Enhanced-FQL($\lambda$) achieves superior sample efficiency and reduced variance compared to n-step fuzzy TD and fuzzy SARSA($\lambda$) baselines, while maintaining substantially lower computational complexity than deep RL alternatives such as DDPG. The framework’s inherent interpretability, combined with its computational efficiency and theoretical convergence guarantees, makes it particularly suitable for safety-critical applications where transparency and resource constraints are essential.
I INTRODUCTION
Designing intelligent agents that can autonomously learn and adapt in high-dimensional, dynamic environments, characterized by pervasive uncertainty and the need for online decision-making, remains a significant challenge in artificial intelligence. To address this core problem, reinforcement learning (RL) has emerged as a powerful framework, with approaches primarily distinguished by their treatment of state and action spaces [1, 2]. A prime example is autonomous robotic navigation, which demands algorithms that are computationally efficient, robust, and capable of online operation [3]. While deep reinforcement learning has demonstrated remarkable success across various domains, its practical deployment encounters several fundamental challenges that limit widespread adoption. A primary concern is the substantial computational requirements of deep neural networks, which often hinder online operation in resource-constrained environments. Additionally, these methods exhibit significant sensitivity to hyperparameter tuning, necessitating extensive experimentation and domain expertise to achieve optimal performance [4]. Perhaps most critically, the black-box nature of deep neural networks results in limited interpretability, thereby restricting their application in safety-critical domains where decision transparency is essential. These limitations collectively motivate the exploration of alternative approaches that can maintain learning capability while offering greater practicality for real-world systems [5].
Building on the RL foundation, value-based methods have demonstrated remarkable effectiveness in discrete action spaces. The Deep Q-Network (DQN) [1] pioneered deep RL by combining Q-learning with neural network function approximation, while Double DQN [6] mitigated overestimation bias through decoupled action selection and evaluation. Dueling DQN [7] further improved learning efficiency by separating value and advantage streams, and Rainbow DQN [8] integrated multiple advancements-including prioritized experience replay and distributional RL-into a unified architecture with superior performance.
In contrast, for continuous control tasks, actor-critic methods have become the dominant paradigm. Deep Deterministic Policy Gradient (DDPG) [9] extended DQN to continuous actions using deterministic policies, while Twin Delayed DDPG (TD3) [10] improved training stability through twin critics and target policy smoothing. More recently, Soft Actor-Critic (SAC) [11] introduced entropy regularization to encourage exploration and improve robustness, making it a strong baseline for complex, continuous domains.
Despite these advances, deep RL methods continue to face fundamental challenges. They typically require extensive hyperparameter tuning and considerable computational resources [12]. Performance is highly sensitive to neural architecture design-such as the number of layers, neurons, and activation functions-yet principled guidelines for selecting these configurations remain elusive [13]. Furthermore, policies represented by deep neural networks often exhibit limited robustness to input noise and environmental variations, and their black-box nature hinders interpretability and formal safety verification in critical applications [14].
These challenges have spurred interest in alternative paradigms that maintain learning capability while offering greater suitability for embedded platforms. Fuzzy systems present a compelling alternative, combining inherent interpretability with built-in robustness to uncertainty through smooth membership functions and substantially lower parameter counts [15, 16]. However, existing fuzzy RL approaches themselves face limitations in scalability and learning efficiency that must be overcome. To address these limitations while preserving interpretability and efficiency, the Fuzzy Q-Learning framework introduced in [17] provides a baseline for integrating fuzzy inference with Bellman equilibrium to achieve robust and interpretable control. While effective in moderate-scale problems, this approach faces scalability limitations when applied to large state-action spaces, leading to slower convergence and lower sample efficiency [18, 19].
In this work, we enhance the Fuzzy Q-Learning structure with a novel fuzzified experience replay mechanism and segmented eligibility traces. This integration enables efficient multi-step credit assignment and better exploitation of past interactions, substantially improving stability and data efficiency. As a result, a more complete and scalable Enhanced Fuzzy Q($\lambda$)-learning approach is introduced, capable of handling complex tasks in continuous state–action space through efficient online learning.
The key contributions of this paper are summarized as: (I) As a novel Fuzzy–RL framework, we introduce the first integration of Fuzzified Bellman Equation (FBE) with a novel fuzzy eligibility traces, TD($\lambda$), and a segmented experience replay, enabling multi-step credit assignment and efficient value propagation across continuous state-action spaces. (II) By replacing deep neural networks with a structured fuzzy rule base, our method eliminates the need for complex architecture design and extensive hyperparameter tuning, drastically reducing the number of learnable parameters while maintaining approximation capability. (III) The smooth interpolation properties of the fuzzy inference system provide built-in robustness to input noise and sensor uncertainties, addressing a key limitation of neural network policies in safety-critical applications. (IV) A contraction mapping-based analysis of the proposed fuzzified Bellman operator is provided, ensuring convergence of the learning process and achieving the suboptimal policy.
II FUNDAMENTAL Q-LEARNING
The fundamentals of standard Q-learning and Fuzzy-QL are explained, here, highlighting their limitations when applied to complex continuous state-action spaces. In a discrete state space $s\in\mathcal{S}$ and a discrete action space $a\in\mathcal{A}$, in Q-learning, based on the Bellman optimality equation, the error of action-value function ($Q(s,a)$) estimation, as a temporal difference (TD), $\delta_{t}$, can be computed. Considering $\gamma\in(0,1)$ as a discount factor and $r_{t}$ as the reward received at step $t$:
$\delta_{t}=r_{t}+\gamma\max_{a^{\prime}}Q(s_{t+1},a^{\prime})-Q(s_{t},a_{t}),$ (1)
Then $Q(s,a)$ will be updated by learning rate $\alpha>0$:
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\delta_{t},$ (2)
While this method is simple and convergent in tabular settings, it is limited to discrete spaces and one-step temporal credit assignment. To address this gap, in [17], Gaussian membership functions are employed to partition continuous state and action spaces:
$\mu_{S_{i}}(s)$ $=\exp\left(-\frac{(s-c_{i}^{S})^{2}}{2\sigma_{S}^{2}}\right),\quad i=1,\ldots,N_{s}$ (3)
$\mu_{A_{j}}(a)$ $=\exp\left(-\frac{(a-c_{j}^{A})^{2}}{2\sigma_{A}^{2}}\right),\quad j=1,\ldots,N_{a}$ (4)
Now, the Fuzzy Q-table, $\widehat{Q}$, contains entries $\widehat{Q}_{i,j}$ associated with each pair of fuzzy membership functions $(\mu_{S_{i}}(s),\mu_{A_{j}}(a))$.
Normalized membership weights, which form a fuzzy distribution over state sets that ensure $\Sigma_{i=1}^{N_{s}}w_{i}(s)=1$, are computed as:
$w_{i}(s)=\frac{\mu_{S_{i}}(s)}{\sum_{k=1}^{N_{s}}\mu_{S_{k}}(s)}$ (5)
The Fuzzified-TD error is then:
$\delta_{i,j}(t)=r_{t}+\gamma\Upsilon(s_{t+1})-\widehat{Q}_{i,j}(t)$ (6)
where the fuzzified next-state value estimation is:
$\Upsilon(s^{\prime})=\sum_{i=1}^{N_{s}}w_{i}(s^{\prime})\max_{j}\widehat{Q}_{i,j}$ (7)
While the FQL-FBE method solved the issue with continuous state-action space, the learning process is long and it is not sample efficient.
III PROPOSED METHOD: Enhanced-FQL($\lambda$)
In this paper we enhanced the Fuzzy–QL with fuzzified eligibility traces and experience replay, to introduce Enhanced-FQL($\lambda$).
III-A Multi-Step Learning with Fuzzified Eligibility Traces
We extend the fuzzy Q-learning with fuzzified eligibility traces for multi-step credit assignment. A Fuzzified Activation Matrix, $\zeta(s,a)$, is introduced as the observed experience. Then, fuzzified eligibility matrix, $0\leq E(t)\leq 1$, is detailed.
$\zeta(s,a)=\mu_{S}(s)\mu_{A}(a)^{\top}$ (8)
$E_{i,j}(t)=\min\{\gamma\lambda E_{i,j}(t-1)+\zeta_{i,j}(s_{t},a_{t}),1\}$ (9)
where $\lambda\in[0,1]$ is the trace-decay parameter. In addition to increasing learning efficiency, this new fuzzified eligibility trace method still maps experiences from their continuous space to a discrete space and re-expresses them in a discrete tabular representation. On the one hand, it does not have the complexity of experiences in continuous space, and on the other hand, it does not have the discrete space limitation of the tabular method in TD($\lambda$). This makes this method suitable for continuous state-action spaces or complex high-dimensional dynamics. Then, the Fuzzy Q-table Bellman update rule combines fuzzified temporal difference with fuzzified eligibility traces:
$\widehat{Q}_{i,j}(t+1)=\widehat{Q}_{i,j}(t)+\alpha E_{i,j}(t)\delta_{i,j}(t)$ (10)
III-B Sample-Efficient Segmented Experience Replay
To enable experience replay while preserving temporal dependencies for eligibility traces, we employ a segment-based replay buffer $\mathcal{D}$:
$\mathcal{D}=\left\{\mathcal{S}_{L}^{(k)}\right\}_{k=1}^{K},\quad\mathcal{S}_{L}=\left\{(s_{\ell},a_{\ell},r_{\ell},s_{\ell+1})\right\}_{\ell=1}^{L}$ (11)
where $\mathcal{D}$ denotes the replay buffer storing $K$ segments, each segment $\mathcal{S}_{L}$ is a sequence of $L$ transitions $(s_{\ell},a_{\ell},r_{\ell},s_{\ell+1})$, with $s_{\ell}$, $a_{\ell}$, $r_{\ell}$, and $s_{\ell+1}$ representing the state, action, reward, and next state at step $\ell$, respectively. Here, a segment refers to a contiguous sequence of state-action-reward transitions $(s_{\ell},a_{\ell},r_{\ell},s_{\ell+1})$ of fixed length $L$.
To maintain temporal consistency in multi-step learning, eligibility traces are reconstructed for each sampled segment. This reconstruction ensures proper credit assignment across consecutive time steps within the segment, which is crucial for the stability and accuracy of the Fuzzy Q($\lambda$)-learning algorithm. This Trace Reconstruction Mechanism is detailed in Algorithm 1.
Algorithm 1 Segment-Based Experience Replay with Trace Reconstruction
1:Input: Replay buffer $\mathcal{D}$, segment length $L$, batch size $B$, $\gamma$, $\lambda$
2:Sample $B$ segments $\{\mathcal{S}_{L}^{(b)}\}_{b=1}^{B}$ from $\mathcal{D}$
3:for each segment $\mathcal{S}_{L}^{(b)}$ do
4:  Initialize $E^{(b)}\leftarrow\mathbf{0}$
5:  for $\ell=1$ to $L$ do
6:   Extract $(s_{\ell},a_{\ell},r_{\ell},s_{\ell+1})$
7:   Compute $\zeta(s_{\ell},a_{\ell})$ via (8)
8:   Update traces: $E^{(b)}$ via (9)
9:   Compute TD error: $\delta_{i,j}$ via (6)
10:  Batch update: $\widehat{Q}\leftarrow\widehat{Q}+\frac{\alpha}{B}\sum_{b=1}^{B}\sum_{\ell=1}^{L}(\delta_{i,j})E^{(b)}$
Experience replay improves sample efficiency by allowing multiple learning updates from each collected experience and decorrelating data samples, resulting in faster and more stable learning with fewer environment interactions.
III-C Action Selection and Exploration
The greedy action is computed through a two-stage process:
$a^{*}(s)=\sum_{i=1}^{N_{s}}p_{i}(s)\cdot c_{j^{*}(i)}^{A}$ (12)
with the policy distribution:
$p_{i}(s)=\frac{\exp\left(w_{i}(s)\max_{j}\widehat{Q}_{i,j}/\beta\right)}{\sum_{k=1}^{N_{s}}\exp\left(w_{k}(s)\max_{j}\widehat{Q}_{k,j}/\beta\right)}$ (13)
and greedy action indices:
$j^{*}(i)=\arg\max_{j}\widehat{Q}_{i,j}$ (14)
An $\epsilon$-greedy exploration strategy is employed during training to balance exploration and exploitation.
III-D Integrated Learning Algorithm
The proposed Enhanced-FQL($\lambda$) algorithm is outlined in Algorithm 2.
Algorithm 2 Fuzzy Q($\lambda$)-Learning with Experience Replay (Enhanced-FQL($\lambda$))
1:Input: Fuzzy partitions, $\alpha$, $\gamma$, $\lambda$, $\epsilon$, $\beta$, $B$, $L$
2:Initialize $\widehat{Q}$, $E\leftarrow 0$, $\mathcal{D}\leftarrow\emptyset$
3:Observe $s_{0}$
4:for $t=0,1,2,\ldots$ do
5:  Compute $\mu_{S}(s_{t})$ via (3)
6:  Compute $w_{i}(s_{t})$ via (5)
7:  Compute $a_{t}$ via $\epsilon$-greedy policy via (12)
8:  Execute $a_{t}$; observe $r_{t}$, $s_{t+1}$
9:  Store transition in current segment
10:  if segment full then
11:   Store segment in $\mathcal{D}$
12:  Compute $\Upsilon(s_{t+1})$ via (7)
13:  Update $E$ via (9)
14:  Compute $\delta_{i,j}$ via (6)
15:  Update $\widehat{Q}$ via (10)
16:  if replay condition then
17:   Execute Algorithm 1
18:  $s_{t}\leftarrow s_{t+1}$
IV Enhanced-FQL($\lambda$) Convergence Study
The convergence and stability of the proposed Enhanced-FQL($\lambda$) method are established under the following standard assumptions.
Assumption 1
The reward function is uniformly bounded, i.e., $|r_{t}|\leq r_{\max}<\infty$. The fuzzy membership functions are bounded within $[0,1]$, and the normalized activation weights $0\leq w_{i}(\cdot)\leq 1$ form a convex combination, satisfying $\sum_{i}w_{i}(\cdot)=1$.
Assumption 2
The environment is modeled as a stationary Markov Decision Process with finite or compact state and action spaces. Under a behavior policy with sufficient exploration (e.g., $\epsilon$-greedy with decaying $\epsilon$), the induced state-action Markov chain is ergodic and admits a unique stationary distribution. Furthermore, the fuzzy rule base experiences persistent excitation, ensuring that all rule pairs $(i,j)$ are sufficiently activated.
Assumption 3
Learning rates ($\alpha_{t}$) satisfy the Robbins–Monro conditions ($\sum_{t=0}^{\infty}\alpha_{t}=\infty$ and $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty$), guaranteeing both sufficient exploration and eventual convergence.
Assumption 4
The discount factor $\gamma$ and trace-decay parameter $\lambda$ lie within $[0,1]$.
Assumption 5
When computing the greedy action via defuzzification, the weight calculation (e.g., using a SoftMax over the per-rule maximum Q-values) ensures non-negative, normalized weights, thereby guaranteeing bounded control outputs and avoiding destabilizing negative weight effects.
Assumption 6
To ensure stability under off-policy learning, Watkins’s Q($\lambda$) is used with trace reset upon exploratory actions.
Definition 1
For any table $Q$, we define a fuzzified optimality operator.
$(\mathcal{T}_{F}Q)_{i,j}\;\triangleq\;\mathbb{E}\!\big[r_{t}\mid(i,j)\big]\;+\;\gamma\,\mathbb{E}\!\big[\Upsilon_{Q}(s_{t+1})\mid(i,j)\big],$ (15)
Boundedness remarks.
(i) Since $E_{i,j}(t)\in[0,1]$, the per-step increment $\alpha_{t}E_{i,j}\delta_{i,j}$ is bounded for bounded $\widehat{Q}$ and rewards. (ii) The greedy action constructed as a convex combination of fixed action centers remains bounded, yielding bounded control inputs.
IV-A Convergence Analysis
Theorem 1
Under Assumptions A1–A6, the sequence $\{\widehat{\mathbf{Q}}_{t}\}$ generated by the Enhanced-FQL($\lambda$) algorithm converges to a fixed suboptimal point $\widehat{\mathbf{Q}}^{\star}$ of the fuzzified Bellman optimality operator $\mathcal{T}_{F}$, i.e., $\lim_{t\to\infty}\widehat{\mathbf{Q}}_{t}=\widehat{\mathbf{Q}}^{\star}$. Consequently, the induced greedy-defuzzified policy converges to the fuzzy-optimal policy over the chosen rule base.
Proof:
The fuzzified Bellman operator $\mathcal{T}_{F}$ is a contraction mapping. In other words, for any two fuzzy Q-functions $\widehat{\mathbf{Q}}_{1}$ and $\widehat{\mathbf{Q}}_{2}$, we have:
$\|\mathcal{T}_{F}\widehat{\mathbf{Q}}_{1}-\mathcal{T}_{F}\widehat{\mathbf{Q}}_{2}\|_{\infty}$
$=\max_{i,j}\left|\mathbb{E}\left[r+\gamma\Upsilon_{1}(s^{\prime})-r-\gamma\Upsilon_{2}(s^{\prime})\right]\right|$
$=\gamma\max_{i,j}\left|\mathbb{E}\left[\Upsilon_{1}(s^{\prime})-\Upsilon_{2}(s^{\prime})\right]\right|$
$\leq\gamma\max_{i,j}\mathbb{E}\left[\left|\Upsilon_{1}(s^{\prime})-\Upsilon_{2}(s^{\prime})\right|\right]$
$\leq\gamma\max_{i,j}\mathbb{E}\left[\sum_{k=1}^{N_{s}}w_{k}(s^{\prime})\left|\max_{l}\widehat{Q}_{1,k,l}-\max_{l}\widehat{Q}_{2,k,l}\right|\right]$
$\leq\gamma\max_{i,j}\mathbb{E}\left[\sum_{k=1}^{N_{s}}w_{k}(s^{\prime})\max_{m}\left|\widehat{Q}_{1,k,m}-\widehat{Q}_{2,k,m}\right|\right]$
$\leq\gamma\max_{i,j}\mathbb{E}\left[\max_{k,m}\left|\widehat{Q}_{1,k,m}-\widehat{Q}_{2,k,m}\right|\right]$
$=\gamma\|\widehat{\mathbf{Q}}_{1}-\widehat{\mathbf{Q}}_{2}\|_{\infty}.$
Since $\gamma\in(0,1)$ by A4, $\mathcal{T}_{F}$ is a $\gamma$-contraction in the sup-norm. By the Banach fixed-point theorem, there exists a fixed suboptimal point $\widehat{\mathbf{Q}}^{\star}$ such that $\mathcal{T}_{F}\widehat{\mathbf{Q}}^{\star}=\widehat{\mathbf{Q}}^{\star}$.
Remark 1
The fuzzy-optimal policy $\pi^{\star}$ is defined through the defuzzified greedy action in Eq. (12)-(14). The convergence of $\widehat{\mathbf{Q}}_{t}$ to $\widehat{\mathbf{Q}}^{\star}$ implies the convergence of $\pi_{t}$ to $\pi^{\star}$ due to the Lipschitz continuity of the SoftMax-based action selection mechanism.
V SIMULATION RESULTS
The proposed Enhanced-FQL($\lambda$) algorithm was evaluated on the classical cart and pole benchmark, a standard continuous control task that presents fundamental challenges in nonlinear stabilization. The environment simulates a pendulum with continuous state and action spaces, where the objective is to swing up and balance the pendulum in the upright position while minimizing control effort. The state space is defined by $s_{t}=[x,\dot{x},\theta,\dot{\theta}]^{\top}$, where $\theta$ represents the pole angle, $\dot{\theta}$ the angular velocity, $x$ the cart position, and $\dot{x}$ the cart linear speed. The action space consists of continuous force values to the cart $a_{t}\in[-2,2]$ N. The reward function is formulated to encourage stabilization with minimal energy consumption:
$r_{t}=-(\theta^{2}+0.1\dot{\theta}^{2}+0.001x^{2}+0.0001\dot{x}^{2}+0.001a_{t}^{2}),$
penalizing state deviation and large control inputs.
For fuzzy approximation, Gaussian membership functions were employed to partition the state space into $N_{s}=\{3,3,7,5\}=315$ fuzzy sets and the action space into $N_{a}=5$ sets. The learning parameters are carefully tuned through experiments, learning rate $\alpha=0.005$, discount factor $\gamma=0.99$, trace decay $\lambda=0.8$, segment length $L=10$, batch size $B=32$, and exploration rate $\varepsilon$ decaying linearly from $0.2$ to $0.05$ over the first 500 episodes to balance exploration and exploitation.
To comprehensively evaluate the performance of the proposed Enhanced-FQL($\lambda$) method, a comparison study is designed.Comparison includes: a multi-step fuzzy Q-learning variant that utilizes $n$-step returns within the same fuzzy rule base, a fuzzy SARSA($\lambda$) approach serving as an on-policy counterpart that employs eligibility traces for policy evaluation, thereby highlighting the impact of the off-policy optimality operator utilized in our method, and DDPG that uses deep neural networks. This comparative implimentation enables a precise analysis of the individual contributions of each algorithmic component to the overall system performance.
Figure 1 illustrates the learning progression over $500$ training episodes, demonstrating the sample efficiency and superior performance of the proposed approach. Enhanced-FQL($\lambda$) achieved consistent convergence (to achieve better Return than -200) within approximately $129$ episodes, significantly outperforming n-step fuzzy Q-learning, fuzzy SARSA($\lambda$), and DDPG, baselines in terms of learning speed and final performance. All comparative evaluations were conducted under identical conditions, including fuzzy partitions, optimizer settings, exploration schedules, and similar random starting of the episode, to ensure a fair comparison.
Figure 1: Learning curves comparing Enhanced-FQL($\lambda$) with baseline methods on Cart&Pole benchmark. The proposed method demonstrates faster convergence and superior performance.
Table I provides a comprehensive quantitative comparison across multiple performance metrics. The proposed method achieved the highest average return. While DDPG delivered competitive performance, it required more computation time per update due to its neural network architecture.
TABLE I: Performance comparison on Cart&Pole environment
Algorithms a b c
n-step FQL 186-186 $0.42$ $388$
Fuzzy SARSA($\lambda$) 197-197 $0.45$ $442$
DDPG 166-166 $0.80$ $379$
Enhanced-FQL($\lambda$) $\mathbf{-159}$ $\mathbf{0.48}$ $\mathbf{129}$
a: Average Return in the last 10% episodes; b: Update Time (ms); Convergence episode
The comparison study validates the effectiveness of integrating fuzzy eligibility traces and segment-based experience replay into FQL–FBE. Three key observations emerge from our analysis:
First, the multi-step credit assignment through fuzzy eligibility traces ($\lambda=0.8$) significantly accelerated learning compared to one-step methods, reducing the sample requirements for convergence by approximately 35% compared to n-step FQL. This demonstrates the importance of temporal credit propagation in continuous control tasks.
Second, the segment-based experience replay mechanism effectively decorrelated training samples while maintaining temporal consistency, resulting in the lowest variance among all evaluated methods. This stability is particularly valuable for real-world applications where training data may be limited or expensive to acquire.
Third, while DDPG achieved respectable performance, Enhanced-FQL($\lambda$) provided a favorable trade-off between performance and computational efficiency. The fuzzy rule base required a few learnable parameters compared to thousands in neural network approaches, making it particularly suitable for resource-constrained embedded systems.
The inherent interpretability of the fuzzy rule base also allows for direct inspection of the learned control strategy, providing valuable insights into the decision–making process—a critical advantage in safety–sensitive applications where neural networks remain black boxes.
Figure 2 compares the speed of reaching a target performance (return = -200). Enhanced-FQL($\lambda$) achieves this with far fewer episodes, demonstrating its superior sample efficiency.
Figure 2: Sample efficiency comparison
VI DISCUSSION
The computed value function $\widehat{Q}^{\star}$ represents the suboptimal fixed point within the chosen fuzzy rule base. While the proposed method involves an inherent bias-approximation trade-off; refining the state and action partitions—that is, increasing $N_{s}$ and $N_{a}$—can reduce approximation bias, it also leads to increased computational complexity. Furthermore, to ensure stability in off-policy control settings, our implementation adopts the Watkins Q($\lambda$) approach, wherein eligibility traces are reset whenever the executed action diverges from the greedy target policy. This mechanism preserves the contraction property of the Bellman optimality operator and facilitates convergent learning behavior.
The bias–variance trade-off and its impact on sample efficiency are fundamental considerations in temporal-difference learning. In $n$-step TD methods, increasing $n$ reduces bias at the cost of higher variance, vice versa. TD($\lambda$) provides a continuous mechanism to balance this trade-off through the trace decay parameter $\lambda$. The proposed Enhanced-FQL($\lambda$) framework inherits this flexible trade-off and further reduces variance by enabling credit assignment across multiple active fuzzy rules via soft generalization. This attribute often leads to improved sample efficiency, particularly in continuous or noisy domains.
The behavior of Enhanced-FQL($\lambda$) can be analyzed through its limit cases with respect to the trace decay parameter $\lambda$. As $\lambda$ approaches zero, the algorithm reduces to one-step fuzzy Q-learning, effectively suppressing eligibility traces and recovering the base FQL-FBE update. Conversely, as $\lambda$ tends toward one, the method approaches Monte Carlo-style credit assignment within the fuzzy basis, propagating returns across entire episodes, subject to episodic finiteness. These limit cases illustrate the continuum of temporal credit assignment strategies encompassed by the proposed framework.
The comparative results reveal three key findings: First, in terms of sample efficiency, Enhanced-FQL($\lambda$) with moderate $\lambda$ values (e.g., 0.5–0.9) achieved target performance levels in fewer episodes compared to $n$-step TD, which can be attributed to continuous credit assignment across activated fuzzy rules. Second, regarding stability, the combination of SoftMax-weighted defuzzification and fuzzy Bellman backups resulted in lower variance in learning curves compared to tabular $n$-step methods, particularly under sensor noise conditions. Third, ablation studies confirmed our design choices by demonstrating performance degradation when using unnormalized value backups ($\Upsilon$), disabling traces ($\lambda=0$), and switching to on-policy SARSA($\lambda$)-style updates.
VII CONCLUSION
This paper has introduced Enhanced-FQL($\lambda$), a novel fuzzy reinforcement learning framework that integrates new fuzzified eligibility traces, segmented experience replay, and a new fuzzy Q-learning architecture. The proposed method addresses key limitations of both deep reinforcement learning and conventional fuzzy RL approaches by offering a computationally efficient and interpretable solution for continuous control tasks. Theoretical analysis has established the convergence guarantee of Enhanced-FQL($\lambda$) under standard assumptions, while empirical results have demonstrated its superior sample efficiency and lower variance compared to n-step fuzzy TD and fuzzy SARSA($\lambda$) baselines. Notably, while achieving competitive performance with deep RL methods like DDPG, our approach maintains significantly lower computational requirements and provides inherent interpretability through its structured fuzzy rule base. The framework’s ability to handle continuous state-action spaces through fuzzy approximation, combined with efficient multi-step credit assignment and memory-aware experience replay, makes it particularly suitable for real-world applications where computational resources are constrained and safety-critical operation requires transparent decision-making.
Building upon these foundations, several promising research directions emerge from this study. First, the integration of prioritized experience replay (PER) could substantially enhance sample efficiency. Segment priority may be defined based on the average or maximum absolute TD error $\|\delta(\ell)\|$ across all activated fuzzy rules. Second, clustering-based experience replay could further optimize memory usage and learning speed and focus on more diverse or informative experiences, reducing redundancy and accelerating policy improvement. Third, extending the framework to multi-agent settings presents a valuable avenue. The interpretable nature of the fuzzy rule base and the sharing of tabular Q data could facilitate coordination and communication among agents. Fourth, exploring online rule adaptation—such as dynamically adjusting the number and placement of fuzzy sets ($N_{s}$, $N_{a}$, $c_{i}$, $\sigma_{i}$) based on incoming data—could enable automatic refinement of the state-action partition, balancing representation power and computational cost over time. Finally, incorporating model-based fuzzy reasoning may allow the agent to simulate and plan over fuzzy state transitions, combining the benefits of model-based reinforcement learning with the interpretability and robustness of fuzzy systems for more efficient long-horizon decision-making.